Show that if the x- and y- intercepts of a line are nonzero numbers a and b, then the equation of the line can be written in the form (x/a) + (y/b) = 1.
The main thing I'm stuck on is what equation to start off with. I've tried point-slope form, and I seem to get there, but I don't understand how to get the 1 into the equation.

Answer by scott8148(6628)

x-intercept is (a,0) , and y-intercept is (0,b)

slope of the line is ___ (0 - b) / (a - 0) = -b/a

using point-slope (with y-int) ___ (y - b) = (-b/a)(x - 0) ___ y - b = -bx/a

dividing by b ___ (y/b) - 1 = -(x/a)

adding (x/a)+1 ___ (y/b) + (x/a) = 1